# inkscapepath=inkscape, inkscapearea=page, clean=true, tex=false

### v1

@@ Conclusion
Patterns of structural discrimination inadvertently introduced into large-scale foundation TTI models are likely to become overshadowed by the biased models and adapters deliberately infused into the digital ecosystem through open-source model-sharing hubs and their users. As shown in this study, on model-sharing hubs like CivitAI biases, particularly those reinforcing hypersexualization and gendered objectification, are amplified and normalized. The rapid rise of NSFW content, the over-representation of young female subjects, and the prioritization of sensational content to drive engagement reflect an exploitative, even abusive dynamic. Additionally, structural discrimination embedded in today’s open-source TTI tools and models have the potential to cause significant downstream harm as they might become widely adopted and even integrated into future consumer applications. Even more alarming is the normalization and incentivization of deepfake model adapters, which fundamentally violate individual dignity, autonomy, and the core democratic principles of self-ownership and consent.
While technological development is not deterministic but contingent, the current trajectory highlights how platform dynamics can exacerbate social pathologies rather than foster equitable access. Yet the contingent nature of technology, shaped by online communities, platform operators, lawmakers, and society as a whole, also creates opportunities for intervention. Model-sharing hubs and social media platforms both have the capacity to implement safeguards that can limit the spread of abusive practices such as deepfake creation and abusive imagery. Strengthening responsible moderation, supporting ethical model development, introducing technical safeguards, and collaborating with NGOs focused on preventing digital violence, are necessary steps to adjust the current trajectory. Rather than reacting after harm has occurred, all stakeholders have the opportunity to innovate proactively and to collectively shape the future direction of model sharing and personalization ecosystems.
@@ Acknowledgments
Figure 10: Addendum to Section 3.1 Typical image and model metadata as obtained via REST API from CivitAI. Some lines have been omitted for clarity. All assets contain unique identifiers assigned by CivitAI, as well as creation dates, usernames etc.
@@
Figure 11: Addendum to Section 3.3.4 Comparison of metadata extracted from downloaded model adapters’ *.safetensors files, showcasing differences between captioning training data with CLIP Interrogator compared to Danbooru-based image captioning systems.
@@
Table 1: Monthly statistics from the Civiverse-40M dataset, covering November 2022 to December 2024. The table reports the total number of generated images, the number classified as NSFW (Not Safe for Work) and non-NSFW, the average number of images generated per day, and the NSFW ratio (the proportion of NSFW images relative to the total image count).
@@
Table 2: Addendum to Section 3.2, MiVOLO Age and Gender Estimations for 2023–2024
@@
Figure 12: Distribution of asset types shared on CivitAI, excluding generated images. Asset types include standalone models (referred to as checkpoints), model adapters (such as LoRA, LoCon, and Textual Inversion), and other resources such as workflows for ComfyUI, pose depth maps, or prompt “Wildcards” shared on the platform. Wildcards refer to placeholder tokens (e.g., early 2000’s flash photography) used to randomly construct prompt variations for diverse image generation. Pose depth maps are structural guides—such as body poses or 3D depth data—used by ControlNets to constrain and compose images.
@@
Figure 13: Danbooru taxonomy of 35.559 unique danbooru tags as extracted from the tag categories and assigned the tag hierarchy from the danbooru.donmai.us tag wiki
Figure 14: Co-occurrence network map of textual training data based on discrete tags, as described in Section 3.4. The data was extracted from the model’s metadata files (*.safetensors) and visualized as a network graph. Each node represents a tag, and edges indicate co-occurrence relationships between tags, with line thickness and proximity reflecting the strength of their association. The tags are color-coded according to the Danbooru taxonomy, providing a visual categorization of semantic clusters and thematic groupings within the dataset. An interactive version of this graph is available at: https://pm-paper-viz.github.io/visualizations/figure_15.html.

### v2

@@ Conclusion
Patterns of structural discrimination inadvertently introduced into large-scale foundation TTI models are likely to become overshadowed by the biased models and adapters deliberately infused into the digital ecosystem through open-source model-sharing hubs and their users. As shown in this study, on model-sharing hubs like CivitAI biases, particularly those reinforcing hypersexualization and gendered objectification, are amplified and normalized. The rapid rise of NSFW content, the over-representation of young female subjects, and the prioritization of sensational content to drive engagement reflect an exploitative, even abusive dynamic. Additionally, structural discrimination embedded in today’s open-source TTI tools and models have the potential to cause significant downstream harm as they might become widely adopted and even integrated into future consumer applications.
Even more alarming is the normalization and incentivization of deepfake model adapters, which fundamentally violate individual dignity, autonomy, and the core democratic principles of self-ownership and consent. While technological development is not deterministic but contingent, the current trajectory highlights how platform dynamics can exacerbate social pathologies rather than foster equitable access. Yet the contingent nature of technology, shaped by online communities, platform operators, lawmakers, and society as a whole, also creates opportunities for intervention. Model-sharing hubs and social media platforms both have the capacity to implement safeguards that can limit the spread of abusive practices such as deepfake creation and abusive imagery. Strengthening responsible moderation, supporting ethical model development, introducing technical safeguards, and collaborating with NGOs focused on preventing digital violence, are necessary steps to adjust the current trajectory. Rather than reacting after harm has occurred, all stakeholders have the opportunity to innovate proactively and to collectively shape the future direction of model sharing and personalization ecosystems.
@@ Acknowledgments
Figure S10: Addendum to Section 3.1 Typical image and model metadata as obtained via REST API from CivitAI. Some lines have been omitted for clarity. All assets contain unique identifiers assigned by CivitAI, as well as creation dates, usernames etc.
@@
Figure S11: Addendum to Section 3.3.4 Comparison of metadata extracted from downloaded model adapters’ *.safetensors files, showcasing differences between captioning training data with CLIP Interrogator compared to Danbooru-based image captioning systems.
@@
Table S1: Monthly statistics from the Civiverse-40M dataset, covering November 2022 to December 2024. The table reports the total number of generated images, the number classified as NSFW (Not Safe for Work) and non-NSFW, the average number of images generated per day, and the NSFW ratio (the proportion of NSFW images relative to the total image count).
@@
Table S2: Addendum to Section 3.2, MiVOLO Age and Gender Estimations for 2023–2024
@@
Figure S12: Distribution of asset types shared on CivitAI, excluding generated images. Asset types include standalone models (referred to as checkpoints), model adapters (such as LoRA, LoCon, and Textual Inversion), and other resources such as workflows for ComfyUI, pose depth maps, or prompt “Wildcards” shared on the platform. Wildcards refer to placeholder tokens (e.g., early 2000’s flash photography) used to randomly construct prompt variations for diverse image generation. Pose depth maps are structural guides—such as body poses or 3D depth data—used by ControlNets to constrain and compose images.
@@
Figure S13: Danbooru taxonomy of 35,559 unique Danbooru tags as extracted from the tag categories and assigned the tag hierarchy from the danbooru.donmai.us tag wiki.
Figure S14: Co-occurrence network map of textual training data based on discrete tags, as described in Section 3.4. The data was extracted from the model’s metadata files (*.safetensors) and visualized as a network graph. Each node represents a tag, and edges indicate co-occurrence relationships between tags, with line thickness and proximity reflecting the strength of their association. The tags are color-coded according to the Danbooru taxonomy, providing a visual categorization of semantic clusters and thematic groupings within the dataset. An interactive version of this graph is available at: https://pm-paper-viz.github.io/visualizations/figure_15.html.